OpenStudy (lila):

Pass attempts A football fan wishes to see how the number of pass attempt, (not completion ) relates to the number of yards gained for quarterback in past NFL season playoff games. The data are shown for five quarter backs. Describe the relationships. Pass atmepts x 116 90 82 108 92 Yards gained y 1001 823 851 873 839

OpenStudy (dumbcow):

is this for a stats class? are you doing a linear regression?

OpenStudy (lila):

yes this is stat

OpenStudy (dumbcow):

ok first find correlation by using mean and std deviation of each x and y r =1/4*sum(((xi-xmean)/sx)((yi-ymean)/sy))

OpenStudy (lila):

i have no clue how that works lolz

OpenStudy (dumbcow):

then find least squares regression line y = b1x+b0 b1 = r(sy/sx) b0 = ymean - b1*xmean

OpenStudy (dumbcow):

lol ok maybe try this then plot the points and see if the line looks linear, quadratic, or exponential

OpenStudy (lila):

lol math is nt my subject i just have to take it